x is a member of A




1
Expert's answer
2021-11-19T12:14:55-0500

Incomplete question:

Let us take an example related to given statement:

Suppose A is bounded and B={x2 | x is a member of A}. Show if sup(A)="\alpha"  then sup(B) = "\alpha"

Solution:

Given that A is bounded sub set of "\mathbb{R}" and "\sup A=\alpha" and "B=\left\{x^{2} / x \in A\right\}"

Since "x \leq \alpha \forall x \in A"

 "\begin{aligned}

&x^{2} \leq \alpha^{2} \\

&x^{2} \leq \alpha^{2} \forall x^{2} \in B

\end{aligned}"

Therefore "\alpha^{2}" is upper bound for B

Let "\beta" be an other upper bound for B

so that "x^{2} \leq \beta \forall x^{2} \in B"

"\Rightarrow x \leq \sqrt{\beta} \forall x \in A"

Therefore "\sqrt{\beta}" is upper bound for A.

Since "\alpha" is suprimum of A, we have "\alpha \leq \sqrt{\beta} \Rightarrow \alpha^{2} \leq \beta" .

Therefore "\alpha^{2}" is least upper bound for B

Hence sup B "=\alpha^{2}"